OpenStudy (marsmcnugget):

What is the solution to the system of equations? y=2x+6 3x-y=6

OpenStudy (marsmcnugget):

@TangerineSkyys

OpenStudy (marsmcnugget):

@petlovers123

OpenStudy (marsmcnugget):

i need to get done fast so i can go see my little sister i havnt seen in months

OpenStudy (petlovers123):

The solutin for y=2x+6 and 3x-y=6 is x = 12 , y = 30 . Equation 1 add equation 2 STEP BY STEP BELOW y + 3x - y = 2x + 6 + 6 3x = 2x + 12 x = 12 y = 2*12 + 6 y = 30

OpenStudy (petlovers123):

hopefully this works :)